Determine the volume of a solution of 1 M HCl that must be added to adjust the pH from 9 to 4 in 100 mL of a 100 mM solution of phosphoric acid. Phosphoric acid has pKs of 2.15, 6.82, and 12.38. For this problem, first estimate the amount of HCl needed based on your knowledge of titration, then accurately determine the correct amount using the Henderson-Hasselbalch equation.
Answer #1

H3PO4(aq) + H2O(l) <==>H2PO4-(aq) + H3O+(aq)

H2PO4-(aq) + H2O(l) <==>HPO42-(aq) + H3O+(aq)

HPO42-(aq)+ H2O(l) <==>PO43-(aq) + H3O+(aq)

we have, PKa1=2.15 ; Pka2 =6.82 ; PKa3 = 12.38

pH = 9.0 , it is after Pka2 , thus based on Pka values we can deduce that :

Principal phosphate species in system : HPO42-(aq)

100 mM = 0.1 M

on adding HCl, we will Have,

HPO42-(aq) + H3O+(aq) <==> H2PO4-(aq) + 2 H2O(l)

we will have ~ 1:1 titration ,

thus V(HCl) ~ 0.1 M *100 mL / 1.0 M ~ 10 mL

pH = pKa1 + log {[H2PO4-]/[H3PO4]}

4.0 = 2.15 + log {[H2PO4-]/[H3PO4]}

log {[H2PO4-]/[H3PO4]} = 1.85

[H2PO4-]/[H3PO4] = 70.8

[H2PO4-] + [H3PO4] = 0.1 M

thus, [H2PO4-] = 0.0986 M; [H3PO4] = 0.0014 M

Thus , amount of HCl needed :

V(HCl) = 1.014*0.1 M *100 mL / 1.0 M = 10.14 mL